i do not understand radical expressions, here is an example -3/2+3/20-3/8
and i do not have the radical mark on my computer

Answer by checkley77(12844)

-3/2+3/20-3/8
you could convert these fractions to decimals thus:
-1.5+.15-.375=-1.725
OR: You could find a common denominator (40) & combine these fractions thus:
(-3*20+3*2-3*5)/40=(-60+6-15)/40=(-75+6)/40=-69/40=-1.725